OpenStudy (anonymous):

To win at LOTTO in one state, one must correctly select 6 numbers from a collection of 62 numbers (1 through 62). The order in which the selection is made does not matter. How many different selections are possible?

OpenStudy (anonymous):

The total number of different 6 number lotto tickets that could be made is:\[62\cdot 62\cdot 62\cdot 62 \cdot 62\cdot 62 = 62^6\]since there are 62 choices for each of the six slots. But order doesnt matter, so we need to divide that by 6! = 720 to get the correct answer.